OpenStudy (anonymous):

What is the length of the midsegment of the trapezoid made by the vertices A(0, 5), B(3, 3), C(5, -2) and D(-1, 2). Show equations and all work that leads to your answer.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

AB has slope -2/3 CD has slope -2/3 They are the parallel sides AB has distance \[\sqrt{9+4}\] CD has distance \[\sqrt{36+16}=\sqrt{52}=2\sqrt{13}\] The distance of the median is \[\frac {AB + CD}{2}\]=\[\frac {3\sqrt{13}}{2}\]